if i am standing on equator, and the earth stops spinning , then around what percent will my weight increase or decrease??? Pls explain!!!

Grade:9

1 Answers

vikas askiitian expert
13 years ago

relation bw accleration due to gravity & angular velocity of earth is given by

geff  = g - w2Rcos@

at equator , @ = 0 ,

R(radius of earth = 6400km &

w = 2pi/T  (T is time of rotation of earth = 24hrs)

after putting these values

geff = 9.76m/s2

 

if earth stops then w becomes 0 &  then

geff = g - 0 = g = 9.8m/s (increased)

change in g = 9.8-9.76 =0.04

percentage change = 0.04/9.76 *100 = 0.4%